OpenStudy (anonymous):

Solve the triangle if possible: a = sqrt(2) c=2sqrt(2) B = 33 degrees.

OpenStudy (anonymous):

First, use the law of cosines to solve for the 3rd side. The law of cosines says: \[b^2=a^2+c^2-2ac\cos\theta\] In this case: \[b^2=2+8-8\cos(33^\circ) = 3.29\] \[b=1.81\] From there, use the law of sines to solve for the other angles. For example: \[\frac{a}{\sin A}=\frac{b}{\sin B}\] \[\sin A=\frac{a\sin B}{b} = 0.42\] \[A = 25.1^\circ\] You can use the same method to find angle C.

OpenStudy (anonymous):

Thank you @srossd